OpenStudy (anonymous):

Two cards are drawn from a standard deck of cards. Find each probability. 1. P(ace, then king) if replacement occurs ace=4/52 king=4/52 (4/52)^2 2. P(heart, then club) if no replacement occurs heart = 13/52 club=13/52 =(13/52)^2 ?

OpenStudy (anonymous):

If no replacement occurs for 2. then surely the p(club) will be 13/51 since there are now only 51 cards left in the pack?

OpenStudy (anonymous):

2) I forgot if no replacement heart = 13/52 club=13/52 (13/52)*(12/51) ?

OpenStudy (anonymous):

1) m'I right?

OpenStudy (anonymous):

No ... if no replacement then (13/52)*(13/51)

OpenStudy (anonymous):

yes, that seems correct.

OpenStudy (anonymous):

you still have 13 clubs but only 51 cards to choose from!

OpenStudy (anonymous):

I'm confusing, which one?

OpenStudy (anonymous):

p(heart then club) = (13/52)*(13/51) asuming no replacement

OpenStudy (anonymous):

ok, than P(ace, then king) ?

OpenStudy (anonymous):

assuming replacement (4/52)*(4/52)

OpenStudy (anonymous):

than 1) I'm correct 2) I just forgot reduce the denominator ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

1. P(ace, then king) if replacement occurs = 4/52 *4/52=1/169 1. P(ace, then king) if replacement occurs 4/52*4/51=4/663

OpenStudy (anonymous):

raheen ... that is valid only if replacement does not occur on the (4/52)*(4/51)

OpenStudy (anonymous):

ty,raheen is correct

OpenStudy (anonymous):

if you take out an ace what is p(king)? 4/51

OpenStudy (anonymous):

P(ace, then king) if replacement occurs than deniminator keep the same P(heart, then club) if no replacement occurs than deniminator reduce to 51

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

think about it in terms of ... if you do not replace the card, how many cards are left in the deck?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

yes so if you want a king and they are all still in the deck then you have a 4/51 chance of picking a king